# Training-Free Exponential Context Extension via Cascading KV Cache

### v2

@@ 4.2 PG19
We measure perplexity on the PG19 (Rae et al., 2019) test set consisting of full-length books. Each book is streamed independently from start to finish without concatenation. We compare against a quadratic flash attention model, as well as BigBird and Streaming LLM, which are also training-free inference adaptations. We use three cache sizes of $16$K, $32$K, and $65$K with a strided prefill of 1K. Since our cache is equivalent to Streaming LLM sequence lengths that are less than the total cache size, we only run each experiment on the subset of books which exceed the given cache size. Flash Attention and Bigbird would exceed the GPU memory capacity, therefore, they are limited to processing books in chunk sizes equivalent to the total cache size. Results are displayed in Table 1. Our method delivers a consistently lower perplexity for all tested cache and model sizes. Additionally, we show examples of how our model behaves during the streaming process in Figure 7. After the cache size is exceeded, our eviction policy begins to differ from Streaming LLM, and our model tends to show lower perplexity due to the increased total token span in our cascading cache. Additional plots for Llama3.1 on all books exceeding 65K length can be seen in Figures 14 and 15.
Table 1: PG19 Perplexity. Across all tested cache sizes, our cascading model maintains lower perplexity than baselines. Flash Attention and Bigbird operate by stepping through the entire sequence with a stride equivalent to the total cache size. They perform attention at each step until reaching the end of the sequence. The Qwen model is excluded from 65K cache size due to having only 32K positional embeddings.
@@
Flash Attn. (strided) Big Bird (strided) Streaming LLM Cascading KV Cache (Ours)
@@
4.3 Passkey Retrieval
We perform passkey retrieval on both Streaming LLM and our Cascading KV Cache. For this experiment, we generate a random 5 digit passkey which is hidden in a random uniform point in the total sequence length. The rest of the text consists of random English words from the dictionary. We perform 20 trials for each insertion location range and sequence length for a total of 600 retrievals. We calculate accuracy for each digit, counting a correct digit prediction if it falls in the proper place in the output sequence. Therefore, a model which outputs random digits would receive an accuracy of 10%. We evaluate total cache sizes of 32K and 65K and sequence lengths with 8 cascades and a strided prefill of 4K. Context lengths start from 32K and double until we reach 1M tokens. Results are shown in Figure 8. For both 32K and 65K cache sizes, Streaming LLM begins to show near random accuracy after the first doubling, while our Cascading KV Cache is still better than random after 4 doublings of the context length.
@@ 4.4 LongBench
We evaluate our method against other linear scaling models on the same subset of tasks as (Xiao et al., 2023) in the LongBench long context understanding benchmark (Bai et al., 2023b). We limit the total cache size of each model to be approximately 1/4 of the original prompt length $L$ of each input using the function $L^{\prime}=2^{\lfloor\log_{2}(L/4)\rfloor}$ and uses a strided prefill of 512. The results are displayed in Figure 9. Averaged over all datasets and tested models, our cascading cache improves performance over the next best model by 12.13%. For tabular results, please see Table 6.
@@ 4.5 Visualization
To visualize the effect of our method on the attention matrices, we reconstruct the full attention matrices of both Streaming LLM and our Cascading KV Cache using Llama3.1 8B on the first 8K tokens of the first book of the PG19 test set. We use a total cache size of 2048 and 4 cascades with a strided prefill of 256. The attention matrices are displayed in Figure 10. Naive sliding window attention Figure 10 (a,c) forms short static barrier where tokens are evicted regardless of their importance. Our method Figure 10 (b,d) maintains those tokens in the context history for a longer time where they may retain influence over future predictions, effectively increasing the available context.
Figure 10: Attention matrix reconstruction for Sink Cache and our Cascading Cache. Both methods result in $\mathcal{O}(n)$ inference time complexity with the same total cache size ($|{\bm{C}}|=2048$).
@@ 4.6 Ablation Study
To study the effect of different parts of our model, we provide three ablation studies including the effect of the strided prefill and token selection using the first book of PG19, and the effect of sparsity induced by the number of cascades. The effect of the strided prefill is shown in Figure 11. We find a decrease in perplexity with an increasing stride size. Intuitively, this comes from the fact that a larger stride provides a larger dense window at the leading edge of the attention matrix as shown in Figure 5. We study the token selection process outlined in Section 3 in Table 3 and find that without the token selection process, our model matches the performance of Streaming LLM, which highlights the importance of selecting higher scoring tokens. Lastly, we study the effect of the number of cascades, and thus overall sparsity, in Figure 12. For this experiment, we use a total cache size of 4K and consider context lengths from 4K which double until 262K. We average the passkey retrieval accuracy over all insertion locations. We find that accuracy steadily increases until the number of cascades exceeds 8 (more than 98% sparse).
@@
Interestingly, the token span in the cache remains a good predictor of accuracy for a moderate number of cascades. Given a token span, we may roughly calculate the expected accuracy in Figure 12(b) by considering the probability that the passkey falls within the span of the tokens. For example, with a token span of 1024, and a context size of 2048, we would expect an accuracy of approximately 50%. We find the token span to be a reliable predictor of accuracy until 4 cascades (73% sparsity).
@@ 5 Limitations & Future Work
As outlined in Equation 4 and visualized in Figure 12, our model’s overall sparsity increases as the number of cascades $N$ grows. This introduces a trade-off, as demonstrated in Figure 12(a), where performance improves up to a point, but diminishes once the number of cascades exceeds eight. Therefore, while our method enables substantial context length extrapolation, this process is not unbounded. Eventually, tokens must be discarded to maintain linear inference complexity, which inherently limits the scope of extrapolation. A promising direction for future work involves addressing the need to discard tokens while preserving linear complexity. One potential solution could involve developing methods for logarithmic complexity searches within the KV cache. By efficiently identifying the top-k relevant key-value pairs, such an approach could eliminate the need for token eviction and allow the model to maintain an overall complexity of $\mathcal{O}(N\log N)$. This would open new avenues for expanding transformer models’ context memory without compromising efficiency.
@@ 6 Conclusion
In this paper, we introduced a novel, training-free method for extending the context memory of streaming LLMs, offering significant improvements without increasing computational complexity. Our approach treats the fixed-size KV cache as a series of cascading sub-caches, allowing for dynamic token retention based on their historical importance. By selectively preserving high-impact tokens and evicting less critical ones, our method effectively extends the context window far beyond the limitations of traditional sliding windows. Our results demonstrate clear performance gains: a 12.13% average improvement on LongBench, a 4.48% boost in Book Summarization tasks, and superior passkey retrieval accuracy at 1M tokens, maintaining a significant edge even after four doublings of the context size. Additionally, our linear prefill strategy eliminates the quadratic complexity of previous approaches, achieving latency reduction by a factor of 6.8 compared to flash attention. These advancements highlight the potential of our method to significantly enhance the efficiency and accuracy of LLMs, making it a practical and impactful solution for both research and real-world applications that require long-context processing.
7 Reproducibility Statement
In order to aid in reproducibility of our experiments, we have provided our code which has been zipped into the supplementary file. We also provide exact pretrained model URL’s which are listed in Table 5. We provide an algorithm for our strided prefill method in Algorithm 1 and a full algorithm for our Cascading KV Cache in Algorithm 2. We have explained the parameters and computation budgets for all experiments in Section 4. As our method is deterministic and requires no stochastic training process, we have omitted error bars in our results.
@@ Appendix A Appendix
  • Appendix B - Extra ablation study on head policy and head reduction function.
@@
Appendix B Head Policy and Head Reduction.
When making a decision for token selection, we may either apply the same homogeneous decision across all heads. Likewise, we may allow the heads to behave independently as illustrated in Figure 13(a). Additionally, as models may make use of Grouped Query Attention (GQA) (Ainslie et al., 2023), the number of attention heads may differ between queries and keys. Therefore, for both cases of homogeneous and independent heads, we need to select a head reduction function which will reduce the head dimension in the attention matrix to 1 (homogeneous heads) or K (key-value heads in GQA). We perform an ablation on the PG19 dataset to explore different options of head reduction functions and head policies in Figure 13(b). We find that in all cases, independent heads outperform homogeneous heads. Among the independent heads, we find that mean and max reductions resulted in similar performance, while a median reduction resulted in slightly worse performance.
@@
Model Method Humanities Soc. Science Other STEM Overall
Llama 3.18B Flash Attention (Full Context) 61.59 76.47 73.09 56.01 67.00
Vanilla (Truncated) 61.23 76.08 73.06 55.79 66.80
Streaming LLM 61.57 76.37 73.19 55.92 66.98
Cascade (Ours) 61.45 76.63 73.25 56.23 67.11
Qwen 27B Flash Attention (Full Context) 63.12 80.40 74.64 64.07 70.99
Vanilla (Truncated) 62.61 80.21 74.64 64.12 70.81
Streaming LLM 63.04 80.47 74.61 63.84 70.86
Cascade (Ours) 62.93 80.40 74.57 63.88 70.87
@@
cascade_cache_buf_array, sink_cache_buf, score_cache, item_to_cache
@@
         $\triangleright$ evict an item from the cache
@@
         if cache_buf.is_empty() then $\triangleright$ eager add to empty cache to avoid naively evicting
@@
              newset_score $\leftarrow$ score_cache.get(newest_item)
              item_score $\leftarrow$ score_cache.get(item_to_cache)
              if item_score ¿ newest_score then
@@
Figure 16: Passkey results for vanilla Llama3.1 as measured on an NVIDIA A6000 GPU with 49GM memory. According to the llama whitepaper (Dubey et al., 2024), it was finetuned for 131K positional embeddings until it achieved 100% accuracy on a passkey retrieval task. Therefore, given enough memory, it should achieve 100 for one more doubling of the context. Our model however, can extend high accuracy numbers past 131K with lower memory usage (see Figure 8).
@@
Table 6: Tabular display of radar plot results from Figure 9. Higher scores are better ($\uparrow$). Total cache sizes are based on the length of the original prompt $L$.

### v3

@@ 4.2 PG19
We measure perplexity on the PG19 test set (full-length books). Each book is streamed independently from start to finish without concatenation. We compare against a quadratic flash attention model, as well as BigBird and Streaming LLM, which are also training-free inference adaptations. We use three cache sizes of $16$K, $32$K, and $65$K with a strided prefill of 1K. Since our cache is equivalent to Streaming LLM sequence lengths that are less than the total cache size, we only run each experiment on the subset of books which exceed the given cache size. Flash Attention 2 and Bigbird would exceed the GPU memory capacity, therefore, they are limited to processing books in chunk sizes equivalent to the total cache size. Results are displayed in Table 1. Our method delivers a consistently lower perplexity for all tested cache and model sizes. Additionally, we show examples of how our model behaves during the streaming process in Figure 7. After the cache size is exceeded, our eviction policy begins to differ from Streaming LLM, and our model tends to show lower perplexity due to the increased total token span in our cascading cache. Additional plots for Llama3.1 on all books exceeding 65K length can be seen in Figures 14 and 15.
Table 1: PG19 Perplexity. Across all tested cache sizes, our cascading model maintains lower perplexity than baselines. Flash Attention 2 and Bigbird operate by stepping through the entire sequence with a stride equivalent to the total cache size. They perform attention at each step until reaching the end of the sequence. The Qwen model is excluded from 65K cache size due to having only 32K positional embeddings.
@@
Flash Attn. 2 (strided) Big Bird (strided) Streaming LLM Cascading KV Cache (Ours)
@@
4.3 Book Summarization
We also evaluate our models on the booksum (Kryściński et al., 2021) book summarization dataset. For this experiment we compare against both linear and quadratic methods Table 2. We find that among linear methods, our models deliver a $4.48\%$ average performance increase over the best linear baseline rouge score (Lin & Hovy, 2003). We find that quadratic methods generally outperform our linear cache, although the performance increases come at a tremendous latency cost for longer input sequences as shown in Figure 6(b).
4.4 Passkey Retrieval
For passkey retrieval, we evaluate Streaming LLM and our Cascading KV Cache, by generating a random 5 digit passkey which is hidden in a random point in the total sequence length. The rest of the text consists of random English words from the dictionary. We perform 20 trials for 5 insertion location ranges and 6 sequence lengths for a total of 600 retrievals. We calculate accuracy for each digit, counting a correct digit prediction if it falls in the proper place in the output sequence. Therefore, a model which outputs random digits would receive an accuracy of 10%. We evaluate total cache sizes of 32K and 65K and sequence lengths with 8 cascades and a strided prefill of 4K. Context lengths start from 32K and double until we reach 1M tokens. Results are shown in Figure 8. For both 32K and 65K cache sizes, Streaming LLM begins to show near random accuracy after the first doubling which exceeds the cache size, while our Cascading KV Cache is still better than random after 4 doublings of the context length.
@@ 4.5 LongBench
We evaluate our method against other linear scaling models on the same subset of tasks as (Xiao et al., 2023) in the LongBench long context understanding benchmark (Bai et al., 2023b). We limit the total cache size of each model to be approximately 1/4 of the original prompt length $L$ of each input using the function $L^{\prime}=2^{\lfloor\log_{2}(L/4)\rfloor}$ and use a strided prefill of 512. Flash attention models truncate text from the middle of the input to achieve the desired cache length. The results are displayed in Figure 9. Averaged over all datasets and tested models, our cascading cache improves performance over the next best model by 12.13%. For tabular results, please see Table 9.
@@ 4.6 Visualization
To visualize the effect of our cache on the attention matrices, we reconstruct the full attention matrices of both Streaming LLM and our Cascading KV Cache using Llama3.1 8B on the first 8K tokens of the first book of the PG19 test set. We use a total cache size of 2048 and 4 cascades with a strided prefill of 256. The attention matrices are displayed in Figure 10. Naive sliding window attention Figure 10 (a,c) forms short static barrier where tokens are evicted regardless of their importance. Our cache Figure 10 (b,d) retains tokens in history for a longer time where have influence over future predictions, effectively increasing the available knowledge.
@@
Figure 10: Attention matrix reconstruction for Streaming LLM’s sink cache and our Cascading Cache. Both methods result in $\mathcal{O}(n)$ inference time complexity with the same total cache size ($|{\bm{C}}|=2048$).
@@ 4.7 Ablation Study
To study the effect of different parts of our model, we provide three ablation studies including the effect of the strided prefill and token selection using the first book of PG19, and the effect of sparsity induced by the number of cascades. The effect of the strided prefill is shown in Figure 11. We find a decrease in perplexity with an increasing stride size. Intuitively, this comes from the fact that a larger stride provides a larger dense window at the leading edge of the attention matrix as shown in Figure 5. We study the token selection process outlined in Section 3 in Table 3 and find that without the token selection process, our model matches the performance of Streaming LLM, which highlights the importance of selecting higher scoring tokens. Lastly, we study the effect of the number of cascades, and thus overall sparsity, in Figure 12. For this experiment, we use a total cache size of 4K and consider context lengths from 4K which double until 262K. We average the passkey retrieval accuracy over all insertion locations. We find that accuracy steadily increases until the number of cascades exceeds 8 (more than 96% window sparsity calculated by $1-|C|/\tilde{S}$).
Interestingly, the token span in the cache remains a good predictor of accuracy for a moderate number of cascades. Given a token span, we may roughly calculate the expected accuracy in Figure 12(b) by considering the probability that the passkey falls within the span of the tokens. For example, with a token span of $\tilde{S}=1024$, and a context size of $2048$, we would expect an accuracy of approximately 50%. We find the token span to be a reliable predictor of accuracy until 4 cascades (Figure 12(c)).
@@ 5 Limitations & Future Work
As described above Equation 4 and visualized in Figure 12, the overall sparsity of our model increases as the number of cascades $N$ increases. This introduces a trade-off, as demonstrated in Figure 12(a), where performance improves up to a point but decreases once the number of cascades exceeds eight. Therefore, while our method enables substantial context length extrapolation, this process is not unbounded. Eventually, tokens must be discarded to maintain linear inference complexity, which inherently limits the scope of extrapolation. A promising direction for future work involves removing the need to discard any tokens while preserving sub-quadratic complexity. One potential solution could involve developing methods for logarithmic complexity searches for query-key similarity within the KV cache. By efficiently identifying the top-k relevant key-value pairs, such an approach could eliminate the need for token eviction and allow the model to maintain an overall complexity of $\mathcal{O}(N\log N)$. This would open new avenues for expanding the context memory of transformer models without compromising efficiency.
@@ 6 Conclusion
In this paper, we introduced a novel, training-free method for extending the context memory of streaming LLMs, offering significant improvements without increasing computational complexity. Our approach treats the fixed-size KV cache as a series of cascading sub-caches, allowing for dynamic token retention based on their historical importance. By selectively preserving high-impact tokens and evicting less critical ones, our method effectively extends the context window far beyond the limitations of traditional sliding windows. Our results demonstrate clear performance gains: a 12.13% average improvement in LongBench, a 4.48% boost in book summarization tasks, and superior passkey retrieval accuracy at 1M tokens, maintaining a significant edge even after four doublings of the context size. Additionally, our linear prefill strategy eliminates the quadratic complexity of previous approaches, achieving latency reduction by a factor of 6.8 compared to flash attention 2. These advancements highlight the potential of our method to significantly improve the efficiency of LLMs, making it a practical and impactful solution for both research and real-world applications that require long-context processing.
Reproducibility Statement
In order to aid in reproducibility of our experiments, we provide our code which is available online [^2] https://github.com/jeffwillette/cascading_kv_cache. We also provide exact pretrained model URL’s which are listed in Table 5. We provide an algorithm for our strided prefill method in Algorithm 1 and a complete algorithm for our Cascading KV Cache in Algorithm 2 and modified flash attention kernnel in Algorithm 3. We have explained the parameters and computation budgets for all experiments in Section 4. As our method is deterministic and does not require a stochastic training or generation process, we have omitted error bars in our results.
Acknowledgments
This work was supported by Institute for Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST)); (No. RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)); (No. RS-2024-00459797, Development of ML compiler framework for on-device AI); (No. RS-2022-II220713, Meta-learning Applicable to Real-world Problems); and (No. RS-2022-00187238, Development of Large Korean Language Model Technology for Efficient Pre-training).
This work was supported by the National Research Foundation of Korea(NRF) grant funded by the Korea government(MSIT) (No. RS-2023-00256259); and (No. RS-2024-00354947).
@@ Appendix A Appendix
  • Appendix B - Extra experimental settings information
  • Appendix C - Extra ablation study on head policy and head reduction function.
@@
Appendix B Experimental Settings
For experiments utilizing 8B model sizes, we use one NVIDIA A6000 (49GB) GPU, and for experiments utilizing 70B model sizes, we utilize 4 NVIDIA A100 GPUs. Quadratic models which use Flash Attention 2 (Dao, 2023) utilize the official cuda kernel, while our method utilizes a modified triton (Tillet et al., 2019) Flash Attention 2 kernel.
Appendix C Head Policy and Head Reduction.
When making a decision for token selection, we may apply the same homogeneous decision across all heads. Likewise, we may allow the heads to behave independently as illustrated in Figure 13(a). Additionally, as models may make use of Grouped Query Attention (GQA) (Ainslie et al., 2023), the number of attention heads may differ between queries and keys. Therefore, for both cases of homogeneous and independent heads, we need to select a head reduction function which will reduce the head dimension in the attention matrix to 1 (homogeneous heads) or K (key-value heads in GQA). We perform an ablation on the PG19 dataset to explore different options of head reduction functions and head policies in Figure 13(b). We find that in all cases, independent heads outperform homogeneous heads. Among the independent heads, we find that mean and max reductions resulted in similar performance, while a median reduction resulted in slightly worse performance. We utilize independent heads and max pooling over attention scores in our experiments.
@@ Appendix B Experimental Settings
Model Method Complexity Humanities Soc. Science Other STEM Overall
Llama 3.18B Flash Attention 2(Full Context) $\mathcal{O}(N^{2})$ 61.59 76.47 73.09 56.01 67.00
Vanilla (Truncated) $\mathcal{O}(N)$ 61.23 76.08 73.06 55.79 66.80
Streaming LLM $\mathcal{O}(N)$ 61.57 76.37 73.19 55.92 66.98
Cascade (Ours) $\mathcal{O}(N)$ 61.45 76.63 73.25 56.23 67.11
Qwen 27B Flash Attention 2 (Full Context) $\mathcal{O}(N^{2})$ 63.12 80.40 74.64 64.07 70.99
Vanilla (Truncated) $\mathcal{O}(N)$ 62.61 80.21 74.64 64.12 70.81
Streaming LLM $\mathcal{O}(N)$ 63.04 80.47 74.61 63.84 70.86
Cascade (Ours) $\mathcal{O}(N)$ 62.93 80.40 74.57 63.88 70.87
@@
cascade_cache_buf_array, sink_cache_buf, item_to_cache
@@
         $\triangleright$ evict an item from the cache and continue
@@
         if not cache_buf.is_full() then $\triangleright$ eager add to unfilled cache to avoid naively evicting
@@
              newest_score $\leftarrow$ cache_buf.get_score(newest_item)
              item_score $\leftarrow$ cache_buf.get_score(item_to_cache)
              if item_score $>$ newest_score then
@@
     update_positional_encoding_indices()
end for
Algorithm 3 Token Selection EMA Accumulation (in the context of Flash Attention 2 kernel)
score_cache, queries, keys, m, EMA beta
for i in chunk(queries) do
     Load $Q_{i}$ from HBM to on-chip SRAM from queries
     for j in chunk(keys) do
         Load $K_{j}$, $V_{j}$ from HBM to on-chip SRAM from keys
         On Chip, Compute $S_{ij}=Q_{i}K_{j}^{\top}$
         On Chip, update $m_{i}$ (update rolling max a la flash attention 2)
         On Chip, update $l_{i}$ (update normalization constant a la flash attention 2)
         On Chip, Compute $\hat{S}_{ij}$ (max-adjusted exponential before softmax normalization)
         On Chip, update $O_{i}$ (update output vector a la flash attention 2)
         On Chip, calculate EMA coeff. $C_{\text{EMA}}$ for $Q_{i}$,
         $\quad\quad\quad C_{\text{EMA}}=\beta^{k}(1-\beta)\forall k\in[$len(queries) - (q_chunk_indices + 1)]]]]
         On Chip, calculate inner loop steps completed $\gamma$ and remaining $\rho$
         Write, Atomic Sum to score_cache += col_sum$((\hat{S}_{ij}/(l_{i}+l_{i}*\frac{\rho}{\gamma}))*C_{\text{EMA}})$
     end for
@@
Refer to captioncache size
Figure 16: Passkey results for vanilla Llama3.1 as measured on an NVIDIA A6000 GPU with 49GM memory and a cache size of 65K. According to the llama whitepaper (Dubey et al., 2024), it was finetuned for 131K positional embeddings until it achieved 100% accuracy on a passkey retrieval task. Therefore, given a big enough cache, it achieves perfect accuracy up to the cache size (or 131K). Our model however, can extend high accuracy numbers well beyond the cache size (see Figure 8).
@@
Model 16K 32K 65K 131K 262K 524K 1M
Minference 5.11 12.84 28.41 60.47 OOM OOM OOM
Cascading KV Cache 9.58 19.71 40.11 80.84 164.47 334.17 665.33
Table 6: Latency (S) as compared to Minference for 1M tokens. This table shows latency throughout all layers, which differs from that shown in Figure 6(b) which shows attention latency for a single layer. Minference goes OOM after 131K on a 49GB GPU due to requiring all tokens in cache for the forward pass. Our model uses a cache size of 16K with a stride of 4K, which are the same settings used in Figure 6(b).
C.1 Additional LongBench/InfiniteBench Baselines
Table 7: Prefill and decoding sparsity for additional models.
Method Sparse Prefill Sparse Decode
MInference
PyramidKV
Cascade (OURS)
In Table 9, we show tabular results for the data displayed in Figure 9. We also add three additional models (MInference (Jiang et al., 2024), PyramidKV (Cai et al., 2024) and Cascade + MInference). The settings between these extra methods contrasts with our setting in various ways (Table 7), as Minference supports a sparse prefill with dense decode and PyramidKV supports a dense prefill with a sparse decode. As our method is the only one which supports both sparse prefill and decode, we truncate the middle of the prompt for MInference and PyramidKV in the same way we do for flash attention in Figure 9.
In addition to this, we have also provded an example of our cache integrated with the official MInference kernel. This allows for using our our linear strided prefill and cascading cache in conjunction with MInference. The key difference in MInference+Cascade is that the leading chunk (stride) of the attention matrix uses MInference instead of dense attention. As MInference is very sparse, this allows for greatly increasing the size of the stride without adding much computation overhead while computing each row of the attention matrix. For this experiment MInference+Cascade used a stride of 100K.
Infinite bench (Zhang et al., 2024a). results can be seen in Table 10 using the same truncation strategy for MInference as outlined above.
C.2 Additional Latency Experiment
In Table 8, we evaluate two additional models (MInference and MInference + Cascade) in terms of latency for a single self attention layer. MInference shows a tendency to scale super-linearly, as each doubling of the context results in a more than doubling of the latency. Because of this, MInference + cascade begins to outperform the latency of plain MInference after sequence lengths of 131K. This decrease in latency is due to the fact that MInference + Cascade uses MInference as the leading chunk in the attention matrix (with a 100K stride) which bounds the total size of the attention matrix of a single attention head at each step to be $\in\mathbb{R}^{100K\times(100K+|C|)}$ where $|C|$ is the total cache size.
Table 8: Additional 1M token latency (S) results. While our cascade model with a stride of 4096 maintains the best overall latency, MInference + cascade also shows linear scaling in addition to added performance benefits shown in Table 9. MInference + cascade begins to outperform plain MInference in terms of latency on sequences over 131K in length. For further discussion of these results, please see Section C.2.
Model 16K 32K 65K 131K 262K 524K 1M
MInference 0.26 0.28 0.78 2.26 6.39 16.84 OOM
MInference + Cascade 0.34 0.68 1.3 3.04 5.77 11.88 24.13
Cascade 0.22 0.46 0.94 1.93 3.81 7.66 15.34
Table 9: Tabular display of LongBench radar plot results from Figure 9. Higher scores are better ($\uparrow$). Total cache sizes are based on the length of the original prompt $L$. For Llama 3.1 8B Instruct we added three additional models (MInference, PyramidKV and Cascade + MInference). For more details regarding these models, please refer to Section C.1.